# The discrete nonlinear Schrödinger equation with linear gain and nonlinear loss: the infinite lattice with nonzero boundary conditions and its finite dimensional approximations

### v1

@@ Abstract
The study of nonlinear Schrödinger-type equations with nonzero boundary conditions define challenging problems both for the continuous (partial differential equation) or the discrete (lattice) counterparts. They are associated with fascinating dynamics emerging by the ubiquitous phenomenon of modulation instability. In this work, we consider the discrete nonlinear Schrödinger equation with linear gain and nonlinear loss. For the infinite lattice supplemented with nonzero boundary conditions which describe solutions decaying on the top of a finite background, we give a rigorous proof that for the corresponding initial-boundary value problem, solutions exist for any initial condition, if and only if, the amplitude of the background has a precise value $A_{*}$ defined by the gain-loss parameters. We argue that this essential property of this infinite lattice can’t be captured by finite lattice approximations of the problem. Commonly, such approximations are defined by lattices with periodic boundary conditions or as it is shown herein, by a modified problem closed with Dirichlet boundary conditions. For the finite dimensional dynamical system defined by the periodic lattice, the dynamics for all initial conditions are captured by a global attractor. Analytical arguments corroborated by numerical simulations show that the global attractor is trivial, defined by a plane wave of amplitude $A_{*}$. Thus, any instability effects or localized phenomena simulated by the finite system can be only transient prior the convergence to this trivial attractor. Aiming to simulate the dynamics of the infinite lattice as accurately as possible, we study the dynamics of localized initial conditions on the constant background and investigate the potential impact of the global asymptotic stability of the background with amplitude $A_{*}$ in the long-time evolution of the system.
@@ I Introduction
Eq. (1.1) is a fundamental nonlinear lattice model incorporating linear and nonlinear gain/loss effects. The parameter $\gamma$ describes linear loss ($\gamma<0$) [or gain ($\gamma>0$)], while $\delta$ describes nonlinear loss ($\delta<0$) [or gain ($\delta>0$)]. The presence of these effects is of crucial importance, particularly in the context of nonlinear optics, where the model  (1.1) may describe the evolution of localized modes in optical waveguides, see [3, 4, 5, 6] (and the references therein). In this context, $\gamma$ describes a linear absorption ($\gamma<0$) [or linear amplification ($\gamma>0$)], while $\delta$ stands for nonlinear amplification ($\delta>0$) [or gain saturation ($\delta<0$)]. The importance of these effects in various discrete and continuous set-ups has been also highlighted in [7, 8, 9, 10, 11, 12] as they may have a prominent role in the potential stability/instability of the localized modes.
In [1] we studied the dynamics of the model  (1.1) in the gain/loss regimes for $\gamma$ and $\delta$, where collapse in finite-time (or blow-up) is manifested. We identified therein, distinct types of collapse and estimates for the blow-up time which in various cases proved to be sharp. We remark that the potential destabilization by finite-time collapse has been proved to be a major characteristic of the dynamical behavior of DNLS models incorporating gain/loss effects. [5, 6].
Herein, we will focus on the study of the dynamics in the case of linear loss $\gamma>0$ and nonlinear gain $\delta<0$. While the loss/loss regime $\gamma<0$, $\delta<0$ is yet of physical significance, it is characterized by the decay of solutions for all initial data. Therefore, our primary concern will be the former regime which is characterized by non-trivial long-time asymptotic behavior.
@@
where $A$ is the amplitude of the constant background. This problem is drastically different from the case of zero boundary conditions and is associated with the emergence of fascinating dynamics and a wide class of localized phenomena which may occur on the top of the finite background $A$, due to the ubiquitous phenomenon of modulation instability (MI).
The second case concerns the system supplemented with periodic boundary conditions. For instance, we will consider an arbitrary number of $N+1$ oscillators which are placed equidistantly at the interval $\Omega=[-L,L]$ of length $2L.$ We denote $k=1/h^{2}$ the discretization parameter, where $h=2L/N$ is the lattice spacing. The spatial coordinate of the oscillators is given then by $x_{n}=-L+nh$, $n=0,1,2,\ldots,N$. Then we supplement Eq. (1.1) with periodic boundary conditions
@@
However, regarding even its local in time solvability, the problem (1.1)-(1.2) is essentially different than its periodic counterpart. We prove that the problem (1.1)-(1.2) admits a unique, at least a local in time solution if and only if $A=A_{*}:=\sqrt{-\gamma/\delta}$. In other words, spatially localized solutions on the top of the finite background $A>0$ exist if and only if $A$ is fixed to the “critical value” $A_{*}$. Apart of its independent interest, this result raises a serious warning concerning the approximations of the infinite lattice problem by the finite lattice, as the periodic one: when $L$ is sufficiently large one may simulate the dynamics of the problem with non-zero boundary conditions (1.2) for arbitrary $A>0$, by triggering initial conditions of the form
@@
In the light of the above result for the infinite lattice, the periodic problem (1.1)-(1.3) deserves much attention. It is shown that the relevant finite dimensional dynamical system posseses a global attractor and its spatially averaged $l^{2}$-norm is uniformly bounded by $A_{*}$. In this paper, to the best of our knowledge, we provide a novel argument that characterizes the global attractor. Specifically, we establish that the global attractor is a time-periodic orbit defined by the unique non-trivial plane wave solution with a constant amplitude denoted as $A_{*}$, a frequency of $\tilde{\omega}$, and a wave number of $q$. This solution satisfies the dispersion relation given by:
@@
What sets our argument apart is that we prove convergence not only in terms of amplitudes, where $A(t)\rightarrow A_{*}$ as $t\rightarrow\infty$ but also through a detailed analysis of the exact frequency function $\Omega(t)$. We determine its behavior in relation to its slant asymptotes, ultimately establishing that $\Omega(t)\rightarrow\Omega_{\infty}(t):=\tilde{\omega}t$. We extend this result to cover the remaining initial conditions through a finite-dimensional representation using their discrete Fourier series. Secondly, we conduct a modulational instability (MI) analysis for the plane wave attractor, demonstrating its potential for modulational instability. By combining the convergence results with the MI analysis, we establish that the MI is transient.
@@
The analysis presented above, particularly in the context of numerical simulations for the periodic lattice, underscores that any localized phenomena arising due to modulational instability (MI) are inherently transient. However, within a finite time frame before reaching ultimate equilibrium, these localized phenomena can be notably intriguing.
For localized initial conditions superimposed on top of $A=A_{*}$, we observe the emergence of waveforms similar to Peregrine solitons as initial events. Subsequently, at later time intervals, spatiotemporal patterns develop due to the modulation instability of the underlying background. These patterns bear a remarkable resemblance to those observed in the dynamics of localized initial data in the integrable Ablowitz-Ladik lattice [13]. They exhibit the universal structure described in [14] during the nonlinear stage of modulational instability. This structure consists of two outer, quiescent sectors corresponding to the supporting background, separated by a wedge-shaped central region characterized by modulated periodic oscillations. The proximity of these dynamics is noteworthy, considering that DNLS (1.1) represents a system that has moved “two steps forward” in breaking the integrability barrier defined by the AL system. DNLS (1.1) is a dissipative perturbation of the non-integrable Hamiltonian DNLS. The stability of the quiescent sectors can be fully justified by the analytical results. Notably, $A=A_{*}$ represents the amplitude of the supporting background, which, as highlighted earlier, is the only amplitude supporting localized solutions in the infinite lattice. Moreover, for the periodic lattice used in the numerical simulations, $A=A_{*}$ corresponds to the amplitude of the globally asymptotically stable background. In the case where $A\neq A_{*}$ where solutions for the infinite lattice do not exist, the numerical simulations for the periodic lattice show that the MI pattern persists on the top, this time of the evolving background. This continues until the amplitude reaches its limiting value $A_{*}$, at which point the entire MI structure disappears and is replaced by the attracting plane wave.
@@
The paper is structured as follows. In Section II, we provide the proof of the existence of solutions for the infinite lattice with nonzero boundary conditions. For the definition and properties of the sequence spaces that define the functional analytic set-up of the problem, we refer to [1] and the references therein. In Section III, we present the results concerning the dynamics of the periodic lattice. Sections IV and V are devoua ted to the numerical results. In particular, in Section IV, we present the results of the numerical studies concerning the transient modulation instability of the plane wave attractor. In Section V, we present the numerical results for localized initial conditions, distinguishing between the case $A=A_{*}$, the only case permitted for the infinite lattice, and the case $A\neq A_{*}$, which can be relevant in the periodic lattice. In this section, we also provide the proof of the estimates for the distance between the solutions of the DNLS (1.1) and the Ablowitz-Ladik periodic lattices, along with the corresponding numerical investigations. We conclude with Section VI, which summarizes our findings.
@@ II The case of the infinite lattice with nonzero boundary conditions
yielding the system in the form
@@
For the local existence of solutions, we may apply the generalized Picard-Lindelöf Theorem [18, Theorem 3.A, pg. 78]. The discrete Laplacian $\Delta_{d}U_{n}=k(U_{n+1}-2U_{n}+U_{n-1})$ is a bounded linear operator $\Delta_{d}:\ell^{p}\rightarrow\ell^{p}$, for any $p\geq 1$, that is, there exists a constant $C>0$, such that
@@
For the operator $F_{1}$, we observe that
$F_{1}(U_{n})$ $=-A^{2}U_{n}-A^{3}+(|U_{n}|^{2}+AU_{n}+A\overline{U}_{n}+A^{2})(U_{n}+A)$
@@
and the embedding (2.11), we may show that $F_{1}:\ell^{p}\rightarrow\ell^{p}$, for any $p\geq 1$, is well defined and is Lipschitz continuous on bounded sets of $\ell^{p}$. That is, for any $U,V\in B_{R}$, with $B_{R}$ a closed ball of $\ell^{p}$ of center 00 and radius $R$, we derive the existence of constants $K_{1}(A,R)$ and $K_{1,1}(A,R)$, such that
$\|F_{1}(U)\|_{\ell^{p}}$ $\leq K_{1}(A,R),$ (2.13)
$\|F_{1}(U)-F_{1}(V)\|_{\ell^{p}}$ $\leq K_{1,1}(A,R)\|U-V\|_{\ell^{p}}.$ (2.14)
Similarly, for the operator $F_{2}$, we have that
$F_{2}(U_{n})=\mathrm{i}\gamma U_{n}+\mathrm{i}\gamma A+\mathrm{i}\delta A^{3}+\mathrm{i}\delta\left\{|U_{n}|^{2}U_{n}+A(U_{n}^{2}+2|U_{n}|^{2})+A^{2}(2U_{n}+\overline{U}_{n}\right\}.$ (2.15)
Assume that $A=A_{*}$. Then, $\mathrm{i}\gamma A+\mathrm{i}\delta A^{3}=0$ and we can show that the operator $F_{2}$ has the properties (2.13)-(2.14). We may proceed by recasting the system (II) in the form of the abstract evolution equation
$\dot{U}=\mathcal{L}(U),\quad\mathcal{L}(U)=\mathrm{i}[-\Delta_{d}(U)-F_{1}(U)+F_{2}(U)],$ (2.16)
@@
For the reverse direction of the proof, we have to show that if a solution of the initial boundary value (1.1)-(1.2)-(1.4) problem exists, then $A=A_{*}$. We argue by contradicion: Assume that a solution of the problem (1.1)-(1.2)-(1.4) exists for all $t\in[0,T]$, for some $T>0$ and that $A\neq A_{*}$. Then, through the same transformations (2.5)-(2.8), a solution $U$ of the problem (II)-(2.7) should exist for all $t\in[0,T]$, and $U\in C^{1}([0,T],\ell^{p})$, which means that $\|\dot{U}(t)\|_{\ell^{p}}<\infty$, for all $t\in[0,T]$. Hence, due to (2.16), $\|\mathcal{L}(U)\|_{\ell^{p}}<\infty$ for all $t\in[0,T]$. This can’t be valid because if $A\neq A_{*}$, the operator $F_{2}$ is not bounded on bounded sets of $\ell^{p}$ as it follows from (2.15). Indeed, for $U$ the solution of the problem which satisfies $\|U(t)\|_{\ell^{p}}\leq R$, for all $t\in[0,T]$, the boundedness of $F_{2}$ in bounded sets of $\ell^{p}$ would imply the counterpart of (2.13),
$\|F_{2}(U(t))\|_{\ell^{p}}\leq K_{2}(A,R),\quad\mbox{for all}\quad t\in[0,T].$ (2.19)
@@
$\mathrm{i}\gamma A+\mathrm{i}\delta A^{3}=-F_{2}(U_{n})+\mathrm{i}\gamma U_{n}+\mathrm{i}\delta\left\{|U_{n}|^{2}U_{n}+A(U_{n}^{2}+2|U_{n}|^{2})+A^{2}(2U_{n}+\overline{U}_{n}\right\},$ (2.20)
@@
In this finite dimensional set-up, it is evident from the definition of the operator $F_{2}$ in (2.15) or (2.20) that Theorem (II.1) is valid for any $A>0$ since the operator $F_{2}:\mathcal{X}^{p}\rightarrow\mathcal{X}^{p}$ is well defined and locally Lipschitz continuous for any $A>0$. Therefore, finite lattice approximations do not capture the essential necessary and sufficient condition on $A$ for the existence of solutions of the infinite lattice which is proved in Theorem II.1.

### v3

@@ Abstract
The study of nonlinear Schrödinger-type equations with nonzero boundary conditions introduces challenging problems both for the continuous (partial differential equation) and the discrete (lattice) counterparts. They are associated with fascinating dynamics emerging by the ubiquitous phenomenon of modulation instability. In this work, we consider the discrete nonlinear Schrödinger equation with linear gain and nonlinear loss. For the infinite lattice supplemented with nonzero boundary conditions, which describe solutions decaying on the top of a finite background, we provide a rigorous proof that for the corresponding initial-boundary value problem, solutions exist for any initial condition, if and only if, the amplitude of the background has a precise value $A_{*}$ defined by the gain-loss parameters. We argue that this essential property of this infinite lattice can’t be captured by finite lattice approximations of the problem. Commonly, such approximations are provided by lattices with periodic boundary conditions or as it is shown herein, by a modified problem closed with Dirichlet boundary conditions. For the finite dimensional dynamical system defined by the periodic lattice, the dynamics for all initial conditions are captured by a global attractor. Analytical arguments corroborated by numerical simulations show that the global attractor is trivial, defined by a plane wave of amplitude $A_{*}$. Thus, any instability effects or localized phenomena simulated by the finite system can be only transient prior the convergence to this trivial attractor. Aiming to simulate the dynamics of the infinite lattice as accurately as possible, we study the dynamics of localized initial conditions on the constant background and investigate the potential impact of the global asymptotic stability of the background with amplitude $A_{*}$ in the long-time evolution of the system.
@@ I Introduction
Eq. (1.1) is a fundamental nonlinear lattice model incorporating linear and nonlinear gain/loss effects. The parameter $\gamma$ describes linear loss ($\gamma<0$) [or gain ($\gamma>0$)], while $\delta$ describes nonlinear loss ($\delta<0$) [or gain ($\delta>0$)]. The important of the presence of these effects is crucial, particularly in the context of nonlinear optics, where the model  (1.1) may describe the evolution of localized modes in optical waveguides, see [3, 4, 5, 6] (and the references therein). In this context, $\gamma$ describes a linear absorption ($\gamma<0$) [or linear amplification ($\gamma>0$)], while $\delta$ stands for nonlinear amplification ($\delta>0$) [or gain saturation ($\delta<0$)]. The importance of these effects in various discrete and continuous set-ups has been also highlighted in [7, 8, 9, 10, 11, 12] as they may have a prominent role in the potential stability/instability of the localized modes.
In [1] we studied the dynamics of the model  (1.1) in the gain/loss regimes for $\gamma$ and $\delta$, where collapse in finite-time (or blow-up) is manifested. We identified therein, distinct types of collapse and estimates for the blow-up time which in various cases proved to be sharp. We remark that the potential destabilization by finite-time collapse has been proved to be a major characteristic of the dynamical behavior of DNLS models incorporating gain/loss effects [5, 6].
Herein, we will focus on the study of the dynamics in the case of linear gain $\gamma>0$ and nonlinear loss $\delta<0$. While the loss/loss regime $\gamma<0$, $\delta<0$ is yet of physical significance, it is characterized by the decay of solutions for all initial data. Therefore, our primary concern will be the former regime which is characterized by non-trivial long-time asymptotic behavior.
@@
where $A$ is the amplitude of the constant background. Note that the specific choice of boundary conditions (1.2) is rather general. The specific choice is imposed by the the unit strength of the nonlinearity in the left-hand side of (1.1), which is selected for simplicity. In the case of a general nonlinearity $f(|u_{n}|^{2})u_{n}$, a consistent problem requires that the right-hand side of the boundary conditions (1.2) must have the form $A\exp[\mathrm{i}f(A^{2})t]$, for a sufficiently smooth function, so that $\lim_{|n|\rightarrow\infty}f(|u_{n}|^{2})=f(\lim_{|n|\rightarrow\infty}|u_{n}|^{2})=f(A^{2})$.
On the one hand, there is a strong motivation from the mathematical point of view for the study of (1.1) with non-zero boundary conditions like (1.2). The problem (1.1)-(1.2) is drastically different from the case of zero boundary conditions and is associated with the emergence of fascinating dynamics and a wide class of localized phenomena which may occur on the top of the finite background $A$, due to the ubiquitous phenomenon of modulation instability (MI). Numerical evidence is provided in [13, Appendix C, pg. 903] for the Hamiltonian non-integrable DNLS ($\gamma=\delta=0$) that the system exhibits the dynamical behavior of modulationally unstable integrable models, like the NLS partial differential equation and the Ablowitz-Ladik (AL) lattice, when supplemented with the boundary conditions (1.2) and the relevant initial conditions of the form of a localized perturbation of a constant background [13, initial data 8(a)-8(c), pg. 891]. The potential persistence of this behavior in the presence of gain and loss effects will be a main question which will be investigated and discussed below. On the other hand, due to the physical relevance of the model (1.1) in the context of optics, it is important to remark that rogue wave-type structures have been observed in remarkable experiments [14], [15]. In the discrete realm, such structures are described by the famous class of discrete rational solutions of the Ablowitz-Ladik lattice [16] which satisfy nonzero boundary conditions like (1.2). The important question of the potential emergence, even robustness of such waveforms in the dynamics of non-integrable DNLS equations initiated new studies as in [17] in the framework of the Salerno lattice and its generalization which may interpolate between the DNLS and the AL system [18]; see also [19], [20] for the existence, stability and dynamics of periodic and quasi-periodic solutions on the top of a finite background for this modified Salerno model. The potential emergence of extreme waveforms in the dynamics of the dissipative DNLS (1.1) will be also investigated in the present work.
The second case concerns the system supplemented with periodic boundary conditions. For instance, we will consider an arbitrary number of $N+1$ oscillators which are placed equidistantly on the interval $\Omega=[-L,L]$ of length $2L.$ We denote $k=1/h^{2}$ the discretization parameter, where $h=2L/N$ is the lattice spacing. The spatial coordinate of the oscillators is given then by $x_{n}=-L+nh$, $n=0,1,2,\ldots,N$. Then we supplement Eq. (1.1) with periodic boundary conditions
@@
However, regarding even its local in time solvability, the problem (1.1)-(1.2) is fundamentally different than its periodic counterpart. We prove that the problem (1.1)-(1.2) admits a unique, at least a local in time solution if and only if $A=A_{*}:=\sqrt{-\gamma/\delta}$. In other words, spatially localized solutions on the top of the finite background $A>0$ exist if and only if $A$ is fixed to the “critical value” $A_{*}$. Apart of its independent interest, this result raises a serious warning concerning the approximations of the infinite lattice problem by the finite lattice, as the periodic one: when $L$ is sufficiently large one may simulate the dynamics of the problem with non-zero boundary conditions (1.2) for arbitrary $A>0$, by triggering initial conditions of the form
@@
In light of the above result for the infinite lattice, the periodic problem (1.1)-(1.3) deserves much attention. It is shown that the relevant finite dimensional dynamical system posseses a global attractor and its spatially averaged $l^{2}$-norm is uniformly bounded by $A_{*}$. In this paper, to the best of our knowledge, we provide a novel argument that characterizes the global attractor. Specifically, we establish that the global attractor is a time-periodic orbit defined by the unique non-trivial plane wave solution with a constant amplitude denoted as $A_{*}$, a frequency of $\tilde{\omega}$, and a wave number of $q$. This solution satisfies the dispersion relation given by:
@@
What sets our argument apart is that we prove convergence not only in terms of amplitudes, where $A(t)\rightarrow A_{*}$ as $t\rightarrow\infty$ but also through a detailed analysis of the exact frequency function $\Omega(t)$. We determine its behavior in relation to its slant asymptotes, ultimately establishing that $\Omega(t)\rightarrow\Omega_{\infty}(t):=\tilde{\omega}t$. We extend this result to cover the remaining initial conditions through a finite-dimensional representation using their discrete Fourier series. Secondly, we conduct a MI analysis for the plane wave attractor, demonstrating its potential for instability. By combining the convergence results with the MI analysis, we establish that this instability is transient.
@@
The analysis presented above, particularly in the context of numerical simulations for the periodic lattice, underscores that any localized phenomena arising due to MI are inherently transient. However, within a finite time frame before reaching ultimate equilibrium, these localized phenomena can be notably intriguing.
For localized initial conditions superimposed on top of $A=A_{*}$, we observe the emergence of waveforms similar to Peregrine solitons as initial events. Subsequently, at later time intervals, spatiotemporal patterns develop due to the MI of the underlying background. These patterns bear a remarkable resemblance to those observed in the dynamics of localized initial data in the integrable Ablowitz-Ladik lattice [13]. They exhibit the universal structure described in [21] during the nonlinear stage of MI. This structure consists of two outer, quiescent sectors corresponding to the supporting background, separated by a wedge-shaped central region characterized by modulated periodic oscillations. The proximity of these dynamics is noteworthy, considering that DNLS (1.1) represents a system that has moved “two steps forward” in breaking the integrability barrier defined by the AL system. DNLS (1.1) is a dissipative perturbation of the non-integrable Hamiltonian DNLS. The stability of the quiescent sectors can be fully justified by the analytical results. Notably, $A=A_{*}$ represents the amplitude of the supporting background, which, as highlighted earlier, is the only amplitude supporting localized solutions in the infinite lattice. Moreover, for the periodic lattice used in the numerical simulations, $A=A_{*}$ corresponds to the amplitude of the globally asymptotically stable background. In the case where $A\neq A_{*}$ where solutions for the infinite lattice do not exist, the numerical simulations for the periodic lattice show that the MI pattern persists on the top, this time of the evolving background. This continues until the amplitude reaches its limiting value $A_{*}$, at which point the entire MI structure disappears and is replaced by the attracting plane wave.
@@
The paper is structured as follows. In Section II, we provide the proof of the existence of solutions for the infinite lattice with nonzero boundary conditions. For the definition and properties of the sequence spaces that define the functional analytic set-up of the problem, we refer to [1] and the references therein. In Section III, we present the results concerning the dynamics of the periodic lattice. Sections IV and V are devoted to the numerical results. In particular, in Section IV, we present the results of the numerical studies concerning the transient MI of the plane wave attractor. In Section V, we present the numerical results for localized initial conditions, distinguishing between the case $A=A_{*}$, the only case permitted for the infinite lattice, and the case $A\neq A_{*}$, which can be relevant in the periodic lattice. In this section, we also provide the proof of the estimates for the distance between the solutions of the DNLS (1.1) and the Ablowitz-Ladik periodic lattices, along with the corresponding numerical investigations. We conclude with Section VI, which summarizes our findings.
@@ II The case of the infinite lattice with nonzero boundary conditions
(in order to make the boundary conditions time independent), yielding the system in the form
@@
For the local existence of solutions, we may apply the generalized Picard-Lindelöf Theorem [25, Theorem 3.A, pg. 78]. The discrete Laplacian
$\Delta_{d}U_{n}=k(U_{n+1}-2U_{n}+U_{n-1}),$
is a bounded linear operator $\Delta_{d}:\ell^{p}\rightarrow\ell^{p}$, for any $p\geq 1$, that is, there exists a constant $C>0$, such that
@@
For the operator $\mathcal{F}_{1}$, we observe that
$\mathcal{F}_{1}(U_{n})$ $=-A^{2}U_{n}-A^{3}+(|U_{n}|^{2}+AU_{n}+A\overline{U}_{n}+A^{2})(U_{n}+A)$
@@
and the embedding (2.11), we may show that $\mathcal{F}_{1}:\ell^{p}\rightarrow\ell^{p}$, for any $p\geq 1$, is well defined and is Lipschitz continuous on bounded sets of $\ell^{p}$. That is, for any $U,V\in B_{R}$, with $B_{R}$ a closed ball of $\ell^{p}$ of center 00 and radius $R$, we derive the existence of constants $K_{1}(A,R)$ and $K_{1,1}(A,R)$, such that
$\|\mathcal{F}_{1}(U)\|_{\ell^{p}}$ $\leq K_{1}(A,R),$ (2.13)
$\|\mathcal{F}_{1}(U)-\mathcal{F}_{1}(V)\|_{\ell^{p}}$ $\leq K_{1,1}(A,R)\|U-V\|_{\ell^{p}}.$ (2.14)
Similarly, for the operator $\mathcal{F}_{2}$, we have that
$\mathcal{F}_{2}(U_{n})=\mathrm{i}\gamma U_{n}+\mathrm{i}\gamma A+\mathrm{i}\delta A^{3}+\mathrm{i}\delta\left\{|U_{n}|^{2}U_{n}+A(U_{n}^{2}+2|U_{n}|^{2})+A^{2}(2U_{n}+\overline{U}_{n}\right\}.$ (2.15)
Assume that $A=A_{*}$. Then, $\mathrm{i}\gamma A+\mathrm{i}\delta A^{3}=0$ and we can show that the operator $\mathcal{F}_{2}$ has the properties (2.13)-(2.14). We may proceed by recasting the system (II) in the form of the abstract evolution equation
$\dot{U}=\mathcal{L}(U),\quad\mathcal{L}(U)=\mathrm{i}[-\Delta_{d}(U)-\mathcal{F}_{1}(U)+\mathcal{F}_{2}(U)],$ (2.16)
@@
For the reverse direction of the proof, we have to show that if a solution of the initial boundary value (1.1)-(1.2)-(1.4) problem exists, then $A=A_{*}$. We argue by contradiction: Assume that a solution of the problem (1.1)-(1.2)-(1.4) exists for all $t\in[0,T]$, for some $T>0$ and that $A\neq A_{*}$. Then, through the same transformations (2.5)-(2.8), a solution $U$ of the problem (II)-(2.7) should exist for all $t\in[0,T]$, and $U\in C^{1}([0,T],\ell^{p})$, which means that $\|\dot{U}(t)\|_{\ell^{p}}<\infty$, for all $t\in[0,T]$. Hence, due to (2.16), $\|\mathcal{L}(U)\|_{\ell^{p}}<\infty$ for all $t\in[0,T]$. This can’t be valid because if $A\neq A_{*}$, the operator $\mathcal{F}_{2}$ is not bounded on bounded sets of $\ell^{p}$ as it follows from (2.15). Indeed, for $U$ the solution of the problem which satisfies $\|U(t)\|_{\ell^{p}}\leq R$, for all $t\in[0,T]$, the boundedness of $\mathcal{F}_{2}$ in bounded sets of $\ell^{p}$ would imply the counterpart of (2.13),
$\|\mathcal{F}_{2}(U(t))\|_{\ell^{p}}\leq K_{2}(A,R),\quad\mbox{for all}\quad t\in[0,T].$ (2.19)
@@
$\mathrm{i}\gamma A+\mathrm{i}\delta A^{3}=-\mathcal{F}_{2}(U_{n})+\mathrm{i}\gamma U_{n}+\mathrm{i}\delta\left\{|U_{n}|^{2}U_{n}+A(U_{n}^{2}+2|U_{n}|^{2})+A^{2}(2U_{n}+\overline{U}_{n}\right\},$ (2.20)
@@
In this finite dimensional set-up, it is evident from the definition of the operators $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ that Theorem II.1 is valid for any $A$ since the operators $\mathcal{F}_{1},\mathcal{F}_{2}:\mathcal{X}^{p}\rightarrow\mathcal{X}^{p}$ are well defined and locally Lipschitz continuous for any $A$. Therefore, finite lattice approximations do not capture the essential necessary and sufficient conditions on $A$, for the existence of solutions of the infinite lattice which is proved in Theorem II.1.